Questions 2-4 use the following data: One beaker c
Answer #1

M = 0.1 M of Zn(NO3)2 and

M = 0.05 M of Al(NO3)3

Q2

the baalnced equation

Al3+ + 3 e− ⇌ Al(s) −1.662

Ni2+ + 2 e− ⇌ Ni(s) −0.25

Eºcell = -0.25 + 1.662 = 1.412

Al will oxidize, since reduciton potential is too low

so

3Ni+2(aq) + 2Al(s) = 2Al+3(aq) + 3Ni(s)

Q3

For this cell

Ecell = Eºcell -0.0592/n * log(Q)

Q = [Al+3]^2 / [Ni+2]^3

Q = (0.05^2)/(0.1^3) = 2.5

n = 6 electrons

Ecell = 1.412 - 0.0592/6*log(2.5) = 1.408073